OpenStudy (anonymous):

multiply (x-5/2)(x+2/5)

OpenStudy (anonymous):

x^2 -(21/10)x - 1

OpenStudy (anonymous):

do i have to collect like terms to get the answer?

OpenStudy (anonymous):

yes collect the "x" terms

OpenStudy (anonymous):

so it would be -2 1/10???

OpenStudy (anonymous):

\[x ^{2} - _{10}^{19}x - 10\]

jhonyy9 (jhonyy9):

(x-5)/2 * (x+2)/5 = ((x-5)(x+2))/10 = (x2 -3x-10)/10